OpenStudy (anonymous):

Solve, using the substitution method. y = 4x + 6 5x – y = 6 A. There are an infinite number of solutions. B. There is no solution. C. (18, 12) D. (12, 54) This would be D right?

OpenStudy (anonymous):

y = 4x + 6 5x – y = 6 => 5x - (4x + 6 ) = 6 => x - 6 = 6 => x = 12 => y = 4*12 + 6 = 6 + 48 = 54... :D hence you are correct again...:D

OpenStudy (anonymous):

:D Thank you!

OpenStudy (anonymous):

welcome :D